# Gravitational wave probes of particle dark matter: a review

### v4

@@ I Introduction
Our knowledge of physics has been incomplete for decades, if not for over a century Bertone and Hooper (2018). Dark energy and dark matter (DM) comprise 69% and 26% of the universe, respectively, and yet we have not been able to directly detect either. The experimental evidence for the existence of DM is overwhelming: stars orbiting the center of the spiral galaxies would have flown out of their orbits if DM did not hold them in place Freeman (1970); Rogstad and Shostak (1972); Whitehurst and Roberts (1972); Roberts and Rots (1973); anisotropies in the cosmic microwave background can result from DM potential wells in the early universe Peebles (1982); Burles and Tytler (1998); DM can seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels and Primack (1982); Davis et al. (1982); Blumenthal et al. (1984); Davis et al. (1985); Navarro et al. (1996); and gravitational lensing of light by galaxies in between us and the light source can be explained by DM inside the galaxy Paczynski (1986); Alcock et al. (1993); Aubourg et al. (1993); Alcock et al. (2000); Lasserre (2000); Tisserand et al. (2007).
@@
While the observational evidence shows that DM exists, it does not explain what it is made of. Can it be a single new particle, such as the axion Peccei and Quinn (1977a, b); Weinberg (1978) or dark photon Galison and Manohar (1984); Holdom (1986); Pierce et al. (2018); Filippi and De Napoli (2020), or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green and Kavanagh (2021)? Or, can particle DM exist alongside PBHs, which may comprise a portion or the totality of DM Hawking (1971); Carr et al. (2010); Dai et al. (2009); Clesse and García-Bellido (2015); Carr et al. (2016); Clesse and García-Bellido (2017); Clesse and Garcia-Bellido (2022); Miller (2024a)?
@@
The numerous hypotheses regarding the origins of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia et al. (2022). Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman et al. (1996). WIMPs have been probed through collider experiments Goodman et al. (2010) and DM/ nucleus interactions by experiments such as CDMS Akerib et al. (2003, 2006); Ahmed et al. (2009), SuperCDMS Agnese et al. (2019), XENON Angle et al. (2008, 2011); Aprile et al. (2012, 2016, 2024), and LZ (LUX-Zeppelin) Malling et al. (2011); Akerib et al. (2013, 2016); Mount et al. (2017); Akerib et al. (2020). Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g. MAGIC Aleksic et al. (2012); Aleksić et al. (2016), VERITAS Holder et al. (2009), HESS Aharonian et al. (2006) and Fermi Hooper and Goodenough (2011), or through cosmic rays, e.g. AMS Aguilar et al. (2007, 2013); Bergstrom et al. (2013) or Fermi-LAT Abdo et al. (2009); Ackermann et al. (2012). Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello and Zurek (2008), and so it is worthwhile to ask whether other models can better explain DM. Thus, the experimental development from the particle physics community has widened to include probing ultralight ($\ll 1$ eV) DM through (1) torsion balance experiments, such as Eöt-Wash Su et al. (1994); Schlamminger et al. (2008) and the MICROSCOPE satellite Touboul et al. (2012); Bergé et al. (2018), whose results can be interpreted as constraints on DM, (2) resonant cavity experiments, such as ADMX Du et al. (2018), (3) the Event Horizon telescope Doeleman et al. (2009); Akiyama et al. (2019), which can measure photons lensed by boson clouds and thus affect BH images Chen et al. (2023), and which can measure changes in the polarization of the electric field when the light passes through the axion background Chen et al. (2022, 2022), and (4) astrometry experiments, such as Gaia Prusti et al. (2016) and the Roman Space Telescope Akeson et al. (2019), that can measure stochastic metric perturbations sourced by ultralight dark matter Kim (2024). In light of the diversity of experiments mentioned, it is also worth asking whether high-precision GW interferometers, such as LIGO, Virgo, KAGRA and GEO600 Affeldt et al. (2014); Dooley et al. (2016); Aasi et al. (2015); Acernese et al. (2015); Aso et al. (2013), can be used to search for DM, as previously reviewed in Bertone et al. (2020); Bertone (2024).
@@
While it has been demonstrated that constraints on axions can be set using current observations of binary neutron star inspirals (GW170817) Hook and Huang (2018); Croon et al. (2018); Zhang et al. (2021) and the Hulse-Taylor binary Kumar Poddar et al. (2020), in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM. CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars. Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska and Bejger (2019); Tenorio et al. (2021); Riles (2023); Piccinni (2022); Miller (2023); Wette (2023). Though the signal model is simple, the unknown sky positions and the uncertain neutron-star physics (e.g. the unknown equation of state, “spin wandering” Akmal et al. (1998); Cutler (2002), changes in the magnetosphere Lyne et al. (2010), etc.) complicate these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan et al. (2004); Dhurandhar et al. (2008); Astone et al. (2014); Suvorova et al. (2016); Bayley et al. (2019). Until recently, such searches have only yielded constraints on the maximum possible deformation of the size of known Abbott et al. (2021a, 2022a); Ashok et al. (2021, 2024); Abac et al. (2025a) and unknown neutron stars Abbott et al. (2022b); Steltner et al. (2023).
@@
    Ultralight bosons (scalar, vector, or tensor): Oscillating fields can couple to standard-model particles, leading to periodic modulations in the interferometer arm lengths, refractive indices, neutron-star moment of inertias, or timings of atomic clocks Khmelnitsky and Rubakov (2014); Pierce et al. (2018); Guo et al. (2019); Grote and Stadnik (2019); Ismail et al. (2023). CW methods are sensitive to these narrowband, quasi-monochromatic signals and can be applied to searches for this kind of DM in ground-based interferometers, pulsar timing arrays, or future space-based antennas. Depending on the particular interferometer, ultralight dark matter masses between $10^{-22},10^{-11}]$ eV can be probed, showing how instruments not even designed to search for DM are able to cover a wide range of masses and thus a variety of DM models.
@@
    Macroscopic dark matter objects: Compact dark objects of masses of $\mathcal{O}(1-10^{9})$ kg passing through or near GW interferometers could induce transient but quasi-periodic disturbances in the test masses Hall et al. (2018); Du et al. (2023). Searches using matched filtering can place constraints on Yukawa-type interactions or on the rate of such transient events. Even though individually, each transient would manifest as short-lived perturbation of the detector, if multiple encounters were to occur over the observation time, the resulting signature would appear as a stochastic background of transients. Potentially, CW methods could be applied to look for the periodicity or stochastic background of such transients. Moreover, networks of interferometers could help discriminate these events from terrestrial noise, providing both spatial and temporal correlations that enhance detection prospects.
@@
    Boson clouds around rotating black holes: Ultralight bosons can grow around black holes via superradiance, accelerated when the Compton wavelength of the boson matches the size of the BH, forming clouds that extract energy and momentum from the BH and emit nearly monochromatic GWs through annihilation or level transitions Arvanitaki and Dubovsky (2011); Arvanitaki et al. (2015a); Brito et al. (2015, 2017). CW searches can target these long-lived signals and test the existence of bosons in the $10^{-13},10^{-11}]$ eV mass range in ground-based GW interferometers, and between $10^{-17},10^{-15}]$ eV in future space-based GW antennas.
@@
    Soliton dark matter: Gravitational interactions can lead to the formation of solitonic cores of ultralight dark matter at the centers of galaxies Schive et al. (2014); Hui et al. (2017). Compact binaries or isolated neutron stars moving through solitons experience additional gravitational potentials that can modify the phase evolution of the GWs as binary systems inspiral, and as neutron stars spin down, respectively. For binaries, this manifests as a long-lived, cumulative dephasing effect during the inspiral Aghaie et al. (2024), while for isolated nonaxisymmetric neutron stars, frequency modulations in the emitted CWs Blas et al. (2025) are induced. In both cases, the effect is coherent over long timescales, making CW search methods particularly well suited to test these scenarios. Depending on the soliton mass and size, signatures would be detectable with pulsar timing array data (probing supermassive binary inspirals crossing galactic cores) or with next-generation terrestrial detectors such as ET and Cosmic Explorer (probing stellar-mass binaries and isolated pulsars embedded in soliton potentials).
@@
    WIMP dark matter: Even though WIMPs are not inherently wave-like, their capture in stars can induce their collapse into BHs Singh et al. (2021); Bhattacharya et al. (2024). CW searches can probe this scenario indirectly, by looking for the long-lived inspirals of BHs space-based GW interferometers. Additionally, null results from CW searches from millisecond pulsars in the Galactic Center can indirectly preference WIMP DM as responsible for the gamma-ray GeV excess emanating from the center of the Milky Way Calore et al. (2019); Miller and Zhao (2023).
@@
    Dark matter spikes around compact objects: Dense DM profiles near BHs inspiraling can generate dynamical friction as the secondary object flows through the DM spike of the first, leading to small but measurable dephasing in these long-lived inspirals Gondolo and Silk (1999); Eda et al. (2013, 2015); Kavanagh et al. (2020); Coogan et al. (2022); Cole et al. (2023a, b); Bertone (2024). While originally such effects were considered around extreme-mass ratio inspirals (EMRIs), intermediate-mass ratio inspirals (IMRIs) and ordinary binaries have been shown to allow the presence of DM spikes, and such dephasing could be measurable in ET, Cosmic Explorer or space-based GW interferometers, when inspiraling binaries will spend hours, days or years in-band. CW methods can thus test the presence and properties of such spikes.
@@
    Atomic dark matter: If DM possesses a composite structure with a dark sector analogue of electromagnetism, it could form bound states or “dark atoms” Feng et al. (2009); Kaplan et al. (2010); Buckley and DiFranzo (2018); Shandera et al. (2018). Such atomic DM can exhibit dissipative dynamics, allowing it to cool and collapse differently from standard cold DM. This can lead to the formation of compact dark objects, modify binary black hole merger populations, or even produce sub-solar mass binaries that would be invisible electromagnetically. While GW searches have constrained only the presence of such atomic DM using matched filtering searches, planetary- or asteroid-mass systems (masses $0.1M_{\odot}$) could also have been formed from atomic DM Fan et al. (2013a, b); Buckley and DiFranzo (2018), and would emit longer-duration GWs. Thus, CW searches can be applied to look for such systems Miller et al. (2022a); Abbott et al. (2022b); Miller et al. (2024).
@@ II.1 Generic features of ultralight dark matter signals
where $v_{0}\simeq 220$ km/s is the virial velocity Smith et al. (2007), $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$cm3 is the DM energy density de Salas and Widmark (2021). The large occupation number $N_{0}$ implies that ultralight DM can be approximated as a plane wave that oscillates at a fixed frequency $f_{0}$, which is given by the DM mass:
@@
In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Pierce et al. (2018).
Despite the fixed frequency implied by Eq. 2, we note that in reality, the DM field is not a perfectly monochromatic wave because the constituent particles have a finite Maxwell-Boltzmann velocity dispersion set by the virial velocity $v_{0}$ of the galactic halo. This spread in velocities causes a small but finite spread in the oscillation frequency, which in turn limits the coherence of the field. As a result, the DM wave can be treated as coherent only over a characteristic time scale $T_{\text{coh}}$, after which the accumulated phase dispersion between different velocity components causes the signal to lose coherence Carney et al. (2021). The wave coherent time $T_{\text{coh}}$ is given by Carney et al. (2021):
@@
Figure 3: Taken from Miller et al. (2021a). We show the strain time series $h(t)$ (left) and the Fourier transform of it (right) of a simulated dark photon DM signal. The structure in the amplitude spectrum density in the right-hand plot arises because we have taken the length of the Fourier transform to exceed that of the coherence time of the signal. This dark photon DM signal is a superposition of $1000$ dark photons traveling with distinct Maxwell-Boltzmann-distributed velocities, which cause small frequency deviations away from the minimal frequency $f_{0}=740.436$ Hz ($m_{A}=3.06\times 10^{-12}$ eV). This plot shows why it is important to choose $T_{\text{FFT}}\sim T_{\text{coh}}$ for the analysis of DM signals. Though we have simulated a dark photon signal, the power spectrum will appear similarly for the other kinds of DM interaction signals. Here, $T_{\text{obs}}\sim 10^{5}$ s, meaning the frequency resolution is $\delta f=10^{-5}$ Hz. The coherence time and length of this signal are: $T_{\text{coh}}=4595$ s and $L_{\text{coh}}=5.28\times 10^{8}$ m; the coupling strength is $\epsilon=3\times 10^{-21}$. The signal is actually simulated for $\sim 233$ days, though we only show the first day of its time evolution.
We will now describe different ways in which DM could interact with particles in the interferometers. We will assume that we are working within $T_{\text{coh}}$, in order to avoid technical difficulties from working with particles of slightly different frequencies. However, we note that in practice, this consideration is necessary when assessing the sensitivity of GW interferometers to DM/standard-model interactions Pierce et al. (2018); Miller et al. (2021a); Vermeulen et al. (2021); Göttel et al. (2024).

### v5

@@ I Introduction
Our knowledge of physics has been incomplete for decades, if not for over a century Bertone:2016nfn. Dark energy and dark matter (DM) comprise 69% and 26% of the universe, respectively, and yet we have not been able to directly detect either. The experimental evidence for the existence of DM is overwhelming: stars orbiting the center of the spiral galaxies would have flown out of their orbits if DM did not hold them in place Freeman:1970mx; rogstad1972gross; whitehurst1972high; roberts1973comparison; anisotropies in the cosmic microwave background can result from DM potential wells in the early universe Peebles:1982ff; Burles:1997fa; DM can seed the formation of large-scale structures visible today, which would have been impossible with ordinary matter alone Pagels:1981ke; davis1982survey; Blumenthal:1984bp; Davis:1985rj; Navarro:1995iw; and gravitational lensing of light by galaxies in between us and the light source can be explained by DM inside the galaxy Paczynski:1985jf; SupernovaCosmologyProject:1993faz; Aubourg:1993wb; MACHO:2000qbb; Lasserre:2000xw; EROS-2:2006ryy.
@@
While the observational evidence shows that DM exists, it does not explain what it is made of. Can it be a single new particle, such as the axion Peccei:1977hh; Peccei:1977ur; Weinberg:1977ma or dark photon Galison:1983pa; Holdom:1985ag; Pierce:2018xmy; Filippi:2020kii, or macroscopic, in the form of PBHs formed within $\sim 1$ second of the Big Bang Green:2020jor? Or, can particle DM exist alongside PBHs, which may comprise a portion or the totality of DM Hawking:1971ei; Carr:2009jm; Dai:2009hx; Clesse:2015wea; Carr:2016drx; Clesse:2016vqa; Clesse:2020ghq; Miller:2024rca?
@@
The numerous hypotheses regarding the origins of DM suggest that the mass of its constituents can vary across hundreds of orders of magnitude Belenchia:2021rfb. Significant time and resources, however, have been spent on searching a somewhat narrow mass range comprising so-called Weakly Interacting Massive Particles (WIMPs) Jungman:1995df. WIMPs have been probed through collider experiments Goodman:2010ku and DM/ nucleus interactions by experiments such as CDMS CDMS:2003yjm; CDMS:2005rss; CDMS:2008uih, SuperCDMS SuperCDMS:2018gro, XENON XENON:2007uwm; XENON10:2011prx; XENON100:2011cza; XENON:2016jmt; XENON:2024wpa, and LZ (LUX-Zeppelin) LZ:2011rhn; LUX:2012kmp; LUX:2016sci; Mount:2017qzi; LZ:2019sgr. Other experiments have tried to detect annihilating DM indirectly through its gamma-ray emission in the Galactic Center, galaxy clusters, or other areas, e.g. MAGIC MAGIC:2011vay; Aleksic:2014poa, VERITAS Holder:2008ux, HESS HESS:2006fka and Fermi Hooper:2010mq, or through cosmic rays, e.g. AMS AMS01:2007rrn; AMS:2013fma; Bergstrom:2013jra or Fermi-LAT Fermi-LAT:2009yfs; Fermi-LAT:2011baq. Despite these efforts, however, WIMPS remain undetected [^1] DAMA claims to see a periodic oscillation of a DM signal, though this result has not yet been reproduced Petriello:2008jj, and so it is worthwhile to ask whether other models can better explain DM. Thus, the experimental development from the particle physics community has widened to include probing ultralight ($\ll 1$ eV) DM through (1) torsion balance experiments, such as Eöt-Wash Su:1994gu; Schlamminger:2007ht and the MICROSCOPE satellite Touboul:2012ui; Berge:2017ovy, whose results can be interpreted as constraints on DM, (2) resonant cavity experiments, such as ADMX Du:2018uak, (3) the Event Horizon telescope Doeleman:2009te; EventHorizonTelescope:2019uob, which can measure photons lensed by boson clouds and thus affect BH images Chen:2022kzv, and which can measure changes in the polarization of the electric field when the light passes through the axion background Chen:2021lvo; Chen:2021lvo, and (4) astrometry experiments, such as Gaia Gaia:2016zol and the Roman Space Telescope Akeson:2019biv, that can measure stochastic metric perturbations sourced by ultralight dark matter Kim:2024xcr. In light of the diversity of experiments mentioned, it is also worth asking whether high-precision GW interferometers, such as LIGO, Virgo, KAGRA and GEO600 Affeldt:2014rza; Dooley:2015fpa; 2015CQGra..32g4001L; 2015CQGra..32b4001A; Aso:2013eba, can be used to search for DM, as previously reviewed in Bertone:2019irm; Bertone:2024rxe.
@@
While it has been demonstrated that constraints on axions can be set using current observations of binary neutron star inspirals (GW170817) Hook:2017psm; Croon:2017zcu; Zhang:2021mks and the Hulse-Taylor binary KumarPoddar:2019jxe, in this review, we would like to explore how continuous gravitational-wave (CWs) can be used as a probe of DM. CWs are quasi-monochromatic, quasi-infinite-duration signals that canonically arise from asymmetrically rotating, lumpy neutron stars. Over the last few decades, much research has focused on developing ways to probe such neutron stars, both those that are known electromagnetically and those that may only be emitting GWs (“gravitars”) Sieniawska:2019hmd; Tenorio:2021wmz; Riles:2022wwz; Piccinni:2022vsd; Miller:2023qyw; Wette:2023dom. Though the signal model is simple, the unknown sky positions and the uncertain neutron-star physics (e.g. the unknown equation of state, “spin wandering” Akmal:1998cf; Cutler:2002, changes in the magnetosphere Lyne:2010ad, etc.) complicate these searches, and thus methods that are robust not only to non-Gaussian noise disturbances but also to theoretical uncertainties had to be developed to look for such systems Krishnan:2004sv; Dhurandhar:2007vb; Astone:2014esa; Suvorova:2016rdc; Bayley:2019bcb. Until recently, such searches have only yielded constraints on the maximum possible deformation of the size of known LIGOScientific:2020lkw; LIGOScientific:2021hvc; Ashok:2021fnj; Ashok:2024fts; LIGOScientific:2025kei and unknown neutron stars KAGRA:2022dwb; Steltner:2023cfk.
@@
    Ultralight bosons (scalar, vector, or tensor): Oscillating fields can couple to standard-model particles, leading to periodic modulations in the interferometer arm lengths, refractive indices, neutron-star moment of inertias, or timings of atomic clocks Khmelnitsky:2013lxt; Pierce:2018xmy; Guo:2019ker; grote2019novel; Ismail:2022ukp. CW methods are sensitive to these narrowband, quasi-monochromatic signals and can be applied to searches for this kind of DM in ground-based interferometers, pulsar timing arrays, or future space-based antennas. Depending on the particular interferometer, ultralight dark matter masses between $10^{-22},10^{-11}]$ eV can be probed, showing how instruments not even designed to search for DM are able to cover a wide range of masses and thus a variety of DM models.
@@
    Macroscopic dark matter objects: Compact dark objects of masses of $\mathcal{O}(1-10^{9})$ kg passing through or near GW interferometers could induce transient but quasi-periodic disturbances in the test masses Hall:2016usm; Du:2023dhk. Searches using matched filtering can place constraints on Yukawa-type interactions or on the rate of such transient events. Even though individually, each transient would manifest as short-lived perturbation of the detector, if multiple encounters were to occur over the observation time, the resulting signature would appear as a stochastic background of transients. Potentially, CW methods could be applied to look for the periodicity or stochastic background of such transients. Moreover, networks of interferometers could help discriminate these events from terrestrial noise, providing both spatial and temporal correlations that enhance detection prospects.
@@
    Boson clouds around rotating black holes: Ultralight bosons can grow around black holes via superradiance, accelerated when the Compton wavelength of the boson matches the size of the BH, forming clouds that extract energy and momentum from the BH and emit nearly monochromatic GWs through annihilation or level transitions Arvanitaki:2010sy; Arvanitaki:2014wva; Brito:2015oca; Brito:2017zvb. CW searches can target these long-lived signals and test the existence of bosons in the $10^{-13},10^{-11}]$ eV mass range in ground-based GW interferometers, and between $10^{-17},10^{-15}]$ eV in future space-based GW antennas.
@@
    Soliton dark matter: Gravitational interactions can lead to the formation of solitonic cores of ultralight dark matter at the centers of galaxies Schive:2014dra; Hui:2016ltb. Compact binaries or isolated neutron stars moving through solitons experience additional gravitational potentials that can modify the phase evolution of the GWs as binary systems inspiral, and as neutron stars spin down, respectively. For binaries, this manifests as a long-lived, cumulative dephasing effect during the inspiral Aghaie:2023lan, while for isolated nonaxisymmetric neutron stars, frequency modulations in the emitted CWs Blas:2024duy are induced. In both cases, the effect is coherent over long timescales, making CW search methods particularly well suited to test these scenarios. Depending on the soliton mass and size, signatures would be detectable with pulsar timing array data (probing supermassive binary inspirals crossing galactic cores) or with next-generation terrestrial detectors such as ET and Cosmic Explorer (probing stellar-mass binaries and isolated pulsars embedded in soliton potentials).
@@
    WIMP dark matter: Even though WIMPs are not inherently wave-like, their capture in stars can induce their collapse into BHs Singh:2020wiq; Bhattacharya:2024pmp. CW searches can probe this scenario indirectly, by looking for the long-lived inspirals of BHs space-based GW interferometers. Additionally, null results from CW searches from millisecond pulsars in the Galactic Center can indirectly preference WIMP DM as responsible for the gamma-ray GeV excess emanating from the center of the Milky Way Calore:2018sbp; Miller:2023qph.
@@
    Dark matter spikes around compact objects: Dense DM profiles near BHs inspiraling can generate dynamical friction as the secondary object flows through the DM spike of the first, leading to small but measurable dephasing in these long-lived inspirals Gondolo:1999ef; Eda:2013gg; Eda:2014kra; Kavanagh:2020cfn; Coogan:2021uqv; Cole:2022ucw; Cole:2022yzw; Bertone:2024rxe. While originally such effects were considered around extreme-mass ratio inspirals (EMRIs), intermediate-mass ratio inspirals (IMRIs) and ordinary binaries have been shown to allow the presence of DM spikes, and such dephasing could be measurable in ET, Cosmic Explorer or space-based GW interferometers, when inspiraling binaries will spend hours, days or years in-band. CW methods can thus test the presence and properties of such spikes.
@@
    Atomic dark matter: If DM possesses a composite structure with a dark sector analogue of electromagnetism, it could form bound states or “dark atoms” Feng:2009mn; Kaplan:2009de; Buckley:2017ttd; Shandera:2018xkn. Such atomic DM can exhibit dissipative dynamics, allowing it to cool and collapse differently from standard cold DM. This can lead to the formation of compact dark objects, modify binary black hole merger populations, or even produce sub-solar mass binaries that would be invisible electromagnetically. While GW searches have constrained only the presence of such atomic DM using matched filtering searches, planetary- or asteroid-mass systems (masses $0.1M_{\odot}$) could also have been formed from atomic DM Fan:2013tia; Fan:2013yva; Buckley:2017ttd, and would emit longer-duration GWs. Thus, CW searches can be applied to look for such systems Miller:2021knj; KAGRA:2022dwb; Miller:2024fpo.
@@ II.1 Generic features of ultralight dark matter signals
where $v_{0}\simeq 220$ km/s is the virial velocity smith2007rave, $m_{\rm DM}$ is the mass of the ultralight dark matter particle, $\lambda_{\rm C}$ is the Compton wavelength of DM, and $\rho_{\text{DM}}=0.3\,\mathrm{GeV}/$cm3 is the DM energy density deSalas:2020hbh. The large occupation number $N_{0}$ implies that ultralight DM can be approximated as a plane wave that oscillates at a fixed frequency $f_{0}$, which is given by the DM mass:
@@
In other words, the observable effect is that the interferometer components will forever oscillate at a fixed frequency since they always sit in the DM field Pierce:2018xmy.
Despite the fixed frequency implied by Eq. 2, we note that in reality, the DM field is not a perfectly monochromatic wave because the constituent particles have a finite Maxwell-Boltzmann velocity dispersion set by the virial velocity $v_{0}$ of the galactic halo. This spread in velocities causes a small but finite spread in the oscillation frequency, which in turn limits the coherence of the field. As a result, the DM wave can be treated as coherent only over a characteristic time scale $T_{\text{coh}}$, after which the accumulated phase dispersion between different velocity components causes the signal to lose coherence Carney:2019cio. The wave coherent time $T_{\text{coh}}$ is given by Carney:2019cio:
@@
Figure 3: Taken from Miller:2020vsl. We show the strain time series $h(t)$ (left) and the Fourier transform of it (right) of a simulated dark photon DM signal. The structure in the amplitude spectrum density in the right-hand plot arises because we have taken the length of the Fourier transform to exceed that of the coherence time of the signal. This dark photon DM signal is a superposition of $1000$ dark photons traveling with distinct Maxwell-Boltzmann-distributed velocities, which cause small frequency deviations away from the minimal frequency $f_{0}=740.436$ Hz ($m_{A}=3.06\times 10^{-12}$ eV). This plot shows why it is important to choose $T_{\text{FFT}}\sim T_{\text{coh}}$ for the analysis of DM signals. Though we have simulated a dark photon signal, the power spectrum will appear similarly for the other kinds of DM interaction signals. Here, $T_{\text{obs}}\sim 10^{5}$ s, meaning the frequency resolution is $\delta f=10^{-5}$ Hz. The coherence time and length of this signal are: $T_{\text{coh}}=4595$ s and $L_{\text{coh}}=5.28\times 10^{8}$ m; the coupling strength is $\epsilon=3\times 10^{-21}$. The signal is actually simulated for $\sim 233$ days, though we only show the first day of its time evolution.
We will now describe different ways in which DM could interact with particles in the interferometers. We will assume that we are working within $T_{\text{coh}}$, in order to avoid technical difficulties from working with particles of slightly different frequencies. However, we note that in practice, this consideration is necessary when assessing the sensitivity of GW interferometers to DM/standard-model interactions Pierce:2018xmy; Miller:2020vsl; Vermeulen:2021epa; Gottel:2024cfj.